OpenStudy (anonymous):

What are the possible number of positive, negative, and complex zeros of f(x) = –2x3 – 5x2 – 6x + 4 ?

OpenStudy (anonymous):

i do not see this being able to be grouped so you shoudl use synthetic division

OpenStudy (anonymous):

What is the highest power of this polynomial???

OpenStudy (anonymous):

oh read it wong...

OpenStudy (anonymous):

Number of roots = Highest Power of the Polynomial...

OpenStudy (anonymous):

Oh im sorry i forgot exponent haha 2x^3 – 5x^2 – 6x + 4 and the highest power is 3

OpenStudy (anonymous):

So roots will also be equal to highest power...

OpenStudy (anonymous):

kk thank you :)

OpenStudy (anonymous):

-6-5(X)^2+6X+4=0 -5(X)^2+6X-2=0 Since the value of D is -4 that is -ve we will have two complex roots. x=(-6+sqrt(D))/(2*(-5)) & x=(-6-sqrt(D))/(2*(-5)) that is x=3/5-i/5 & x=3/5+i/5

OpenStudy (anonymous):

two complex roots, no positive or negative zeros